OpenStudy (anonymous):

A tree is supported by a wire anchored in the ground 12 feet from it's base. The wire is 5 feet longer than the height that it reaches on the tree. Find the length of the wire. The wire is ??? feet long

OpenStudy (anonymous):

@Agent47

OpenStudy (anonymous):

@im_thompson5910

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (agent47):



OpenStudy (agent47):

use the pythagorean theorem.

OpenStudy (anonymous):

what is that

jimthompson5910 (jim_thompson5910):

solve h^2 + 12^2 = (h+5)^2 for h

OpenStudy (anonymous):

i got h^2+10h+25?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

(h+5)^2 becomes h^2+10h+25 so you're correct in that aspect

jimthompson5910 (jim_thompson5910):

so h^2 + 12^2 = (h+5)^2 turns into h^2 + 144 = h^2 + 10h + 25

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so is that my answer

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

no, you want to find the value of h once you find h, you can find h+5

OpenStudy (anonymous):

ok so how do i do that

jimthompson5910 (jim_thompson5910):

What do you see canceling in that last equation above?

OpenStudy (anonymous):

h^2?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

ok so what's next

jimthompson5910 (jim_thompson5910):

so we have 144 = 10h + 25 solve that for h

OpenStudy (anonymous):

would it be 10h+169